OpenStudy (anonymous):

How to evaluate -> (5^-4 - 5^-6)/(5^-3 +5^-5)

OpenStudy (anonymous):

bet you got the idea now. write it without the negative exponents. then clear the fractions

OpenStudy (anonymous):

Okay so:

OpenStudy (anonymous):

\[\frac{5^{-4} - 5^{-6}}{5^{-3} +5^{-5}}\] muttiply top and bottom by \[5^6\]

OpenStudy (anonymous):

I got (1/5^4 - 1/5^6)/ (1/5^3 +1/5^5)

OpenStudy (anonymous):

wait. How come I have to multiply top and bottom by 5^6?

OpenStudy (anonymous):

that will clear the fractions. \[5^{-4}\times 5^6=5^2\]

OpenStudy (anonymous):

if you want you can write this in fraction notation. then you will see that you should multiply top and bottom by \[5^6\]

OpenStudy (anonymous):

i will write it out if you like

OpenStudy (anonymous):

please do, thanks

OpenStudy (anonymous):

\[\frac{5^{-4}-5^{-6}}{5^{-3}+5^{-5}}\times \frac{5^6}{5^6}\]

OpenStudy (anonymous):

\[\frac{5^2-1}{5^3+5}\] is the second line clear?

OpenStudy (anonymous):

pardon?

OpenStudy (anonymous):

i added the exponents, and i chose \[5^6\] so that they would all be positive

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

because i want to evaluate the numbers, i need the exponents to be positive

OpenStudy (anonymous):

mmhm

OpenStudy (anonymous):

now i can see what the numbers are. numerator is 25+1=26

OpenStudy (anonymous):

denominator is 125+5=130

OpenStudy (anonymous):

Isn't the numerator 25-1?

OpenStudy (anonymous):

oops yes

OpenStudy (anonymous):

Okay, that's great, thanks :) I was making it way more complicated thn it needed to be

OpenStudy (anonymous):

i would like to say that was a "typo" but i guess it was a mistake. in any case you have the answer, it is \[\frac{24}{130}=\frac{12}{65}\]

OpenStudy (anonymous):

good! yw

OpenStudy (anonymous):

i like ur pic